During the first half of a basketball game, Emily missed all five of her field goal attempts. During the second half, she hit 75% of her 16 attemplts. What was her goal shooting percentage for the game?
Answer by rajagopalan(174)

First half total goals 0,attemt 5
Second half total 16x3/4=12 goals, 16 attemts
so total goals 12
total attemps 21
percentage =12/21=4/7=57.1%